Evaluate



a) sin (alpha), given that sec (alpha)= -3, and that alpha is an obtuse angle.


b) cot(beta), given that tan(beta)=1/(√2) and that beta is an acute angle.

1
Expert's answer
2022-05-06T11:15:41-0400

a) "90\degree<\alpha<180\degree"


"\cos \alpha=\dfrac{1}{\sec \alpha}=-\dfrac{1}{3}"

"\sin \alpha=\sqrt{1-\cos^2 \alpha}=\sqrt{1-(-\dfrac{1}{3})^2}=\dfrac{2\sqrt{2}}{3}"

b) "0\degree<\alpha<90\degree"


"\cot \beta=\dfrac{1}{\tan \beta}=\sqrt{2}"